Al + Fe3O4 → Fe + Al2O3

Determine the masses of all the substances present after the reaction if  187 g of Al and  791 g of Fe3O4 react to completion. Some amounts may be zero (0).



1
Expert's answer
2021-09-23T02:37:10-0400

8Al + 3Fe3O4 → 9Fe + 4Al2O3

M(Al) = 26.98 g/mol

"n(Al) = \frac{187}{26.98} = 6.931 \;mol"

M(Fe3O4) = 231.53 g/mol

n(Fe3O4) "= \frac{791}{231.53} = 3.416 \;mol"

According to the reaction equation for each 3 mol of Fe3O4 we need 8 mol of Al, but we have only 3.416 mol of Fe3O4 for 3.416 mol of Al. So, Fe3O4 is a limiting reactant.

According to reaction:

From one mole of Fe3O4 we can obtain 3 mol of Fe, so

n(Fe) =3n(Fe3O4) "= 3 \times 3.416 = 10.249 \;mol"

M(Fe) = 55.84 g/mol

m(Fe) "= 55.84 \times 10.249 = 572.31 \;g"

From 3 mol of Fe3O4 we can obtain 4 mol of Al2O3, so

n(Al2O3) "= \frac{4}{3}n(Fe_3O_4) = \frac{4}{3} \times 3.416 = 4.554 \;mol"

M(Al2O3) = 101.96 g/mol

m(Al2O3) "= 101.96 \times 4.554 = 464.39 \;g"